Etch Clean Graphics uses one cleanser that is 25% acid and a second that is 50% acid. How many liters of each should be mixed to get 10 L of a solution that is 40% acid?
Answer by jorel1380(3719)

10x.40=4
m+n=10
.25m+.50n=4
25m+50n=400
25m+25n=250
25n=150
n=6
m=4
You need 4 liters of 25% acid and 6 liters of 50% acid
(4x.25)+6(.5)=1+3=4=.4(10)